OpenStudy (jamaljunior):

What is the value of g(−3) when g(x)=2x−2 ?

OpenStudy (will.h):

Substitute x = -3 in the function

OpenStudy (jamaljunior):

how?

OpenStudy (will.h):

G(x) = 2x - 2 When x = -3 G(-3) = 2(-3) - 2 Simplify

OpenStudy (jamaljunior):

G-3 = -6 - 2?

OpenStudy (jamaljunior):

because 2x -3

OpenStudy (will.h):

You can't simplify G(-3) just focus of 2(-3) -2 What would the result be

OpenStudy (will.h):

The total result

OpenStudy (will.h):

G(-3) =??

OpenStudy (jamaljunior):

-8

OpenStudy (will.h):

Correct That's it

OpenStudy (jamaljunior):

-6 - 2?

OpenStudy (jamaljunior):

Thanks so much!

OpenStudy (will.h):

Sure